OpenStudy (anonymous):

Find the volume of a cube with the given side length. 10 units A. 10,000 cubic units B. 1000 cubic units C. 100 cubic units D. 30 cubic units

OpenStudy (anonymous):

10^3 = 1000 You should be able to do this by yourself, considering that I just posted to you how to solve these types of problems .

OpenStudy (anonymous):

okay so you just times them bye three?

OpenStudy (anonymous):

yes 10^3=10*10*10

OpenStudy (anonymous):

okay so like if its 11 you times 11 by three and you will get 33

OpenStudy (anonymous):

No you raise it to the third power, so it would be 11*11*11

OpenStudy (anonymous):

so that would be 1331

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks i have one that i dont get but yeah